# Feedback Linearisation with State Constraints

### v1

@@ Abstract
[Abstract]Feedback Linearisation (FBL) is a widely used technique that applies feedback laws to transform input-affine nonlinear dynamical systems into linear dynamical systems, allowing for the use of linear controller design methods such as pole placement. However, for problems with state constraints, controlling the linear system induced by FBL can be more challenging than controlling the original system. This is because simple state constraints in the original nonlinear system become complex nonlinear constraints in the FBL induced linearised system, thereby diminishing the advantages of linearisation. To avoid increasing the complexity of state constraints under FBL, this paper introduces a method to first augment system dynamics to capture state constraints before applying FBL. We show that our proposed augmentation method leads to ill-defined relative degrees at state constraint boundaries. However, we show that ill-defined relative degrees can be overcome by using a switching FBL controller. Numerical experiments illustrate the capabilities of this method for handling state constraints within the FBL framework.
@@ 1 Introduction
Nonlinear systems pervade every facet of our world, from turbulent flows [1] to the sophisticated operations of human-made cybernetic infrastructure [2]. Unlike linear systems, nonlinear systems do not obey the solution superposition property leading to phenomena such as non-global stability, limit cycles and sensitivity to small initial changes (chaos). These properties make the control of nonlinear systems significantly more challenging than that of linear systems. Nevertheless, there exist several techniques that approximate or transform challenging nonlinear problems into simpler linear problems, such as Jacobean/Taylor linearisation, Koopman operators [3], Carleman linearisation [4] and Feedback Linearisation. FBL is a fairly mature topic based on the early works [5, 6] from the seventies and has been extensively treated in textbooks [7].
@@
The contribution of this paper is to propose a general FBL method capable of handling state constraints without the need for any online optimisation. Our method is summarised in the flowchart given in Fig. 1, follows three steps: 1) Augment the dynamics to capture state constraints and modify with integral control structure, 2) Apply FBL to the integral augmented system to derive an FBL controller, 3) Implement the FBL controller where augmented state dependencies are removed via algebraic substitution.
@@
Extending the state space by augmenting the system dynamics has previously been used to solve discrete-time problems with non-additively separable cost functions [29, 30]. Our work focuses on tracking problems in continuous time rather than minimising general cost functions, using state augmentation to encapsulate state constraints. We follow and extend the approach of [31] by introducing a slack variable to represent the square root of the proximity to the state constraint boundary. Analogous to the first step of FBL, we differentiate this equation to establish a relationship between the slack variable and the system input. Solving for the input, we obtain a new higher dimensional system where the virtual input corresponds to the highest derivative of the slack variable. We show that the augmented system will not violate state constraints for any integrable input. We overcome the limitation of [31], that the number of input dimensions must match the number of constraints required, by proposing a sequential approach where we iteratively augment the state space to capture constraints. We also differ from [31] by imposing an integral controller structure to ensure any input used is integrable and hence cannot violate the state constraints. Moreover, unlike in [31], which uses HJB equations to solve for the controller, we synthesise a closed-loop controller using FBL and analyse the associated challenges that arise from this approach.
@@
Unfortunately, it is non-trivial to apply FBL out of the box to the resulting augmented system, since we show that the augmented system has an ill-defined relative degree, changing its value along the state constraint boundary. Therefore, in order to apply FBL, we use the switching controller from [32] [33] that was designed to handle cases of ill-defined relative degree. Finally, we synthesise a state constraint feasible closed-loop controller for the original system coupled with integral states by removing the augmented states such that the controller only depends on the states of the original system and the induced integral states.
@@
While our underlying methodology is applicable to MIMO systems, the implementation involving a switching controller is restricted to Single-Input Single-Output (SISO) systems. Accordingly, Sections 2 to 4, which extend the state augmentation approach of [31] and analyse the complications that arise when combining this method with FBL, are presented in the general MIMO setting. From Section 5 onwards, however, we focus exclusively on SISO systems, where the practical implementation of the FBL controller, particularly the switching logic, is more tractable.
@@
We denote the set of $k$-differentiable functions, $f:X\to Y$ by $C^{k}(X,Y)$, with $C^{\infty}(X,Y)$ denoting the set of infinitely differentiable functions. Along the same lines, we denote $L^{1}(X,Y)$ as the set of Lebesgue integrable functions. Given $F\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{r})$ and $x\in C^{1}([0,\infty),\mathbb{R}^{n})$ we define $\frac{dF(t,x(t))}{dt}=\frac{\partial F(t,x(t))}{\partial x}\frac{\partial x(t)}{\partial t}+\frac{\partial F(t,x(t))}{\partial t}$, where $\frac{\partial F(t,x)}{\partial x}=[\frac{\partial F(t,x)}{\partial x_{1}},\dots,\frac{\partial F(t,x)}{\partial x_{n}}]\in\mathbb{R}^{1\times n}$ is a row vector. We define the 0-th derivative as $\frac{d^{0}}{dt^{0}}F(t,x)=F(t,x)$ and $\frac{\partial^{0}}{\partial t^{0}}F(t,x)=F(t,x)$. Consider $\psi\in C^{i}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{m})$, $\psi^{(i)}(t,x(t))$ denotes the i-th order time derivative of $\psi(t,x(t))$. Given $f\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{n})$ and $g\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{n\times m})$ we denote the Lie derivatives of $\psi\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{m})$, by $L_{f}^{j}\psi$ and $L_{g}L_{f}^{j}\psi$ that are are given in Eqs. (1).
$L_{f}^{0}\psi(t,x)=\psi(t,x),$
$L_{f}^{j}\psi(t,x)=\frac{\partial L_{f}^{j-1}\psi(t,x)}{\partial x}f(t,x)+\frac{\partial L_{f}^{j-1}\psi(t,x)}{\partial t}\text{ for }j\in\mathbb{N}$
$L_{g}L_{f}^{j}\psi(t,x)=\frac{\partial L_{f}^{j}\psi(t,x)}{\partial x}g(t,x)\text{ for }j\in\mathbb{N}\cup\{0\}.$ (2)
This definition of Lie derivative is useful when dealing with time-varying systems, see [34] for more details.
@@ 3 State Augmentation: From
The slack variable, $Z$, depends explicitly on the state of the system through Eq. (8) and thus implicitly depends on the input of the system. By computing the explicit expression that relates the system input to the slack variable, we may substitute the input variable for the slack variable, treating the slack variable as a pseudo-input variable of the transformed system. Then, the transformed system will remain state-feasible for any value taken by the pseudo-input, since by the quadratic nature of the slack variable it follows from Eq. (8) that $\phi(t,x(t))\geq 0$ for any $t\in[t_{0},\infty)$. In the work of [31], the time derivatives of Eq. (8) were not analytically derived. Next, we extend this work by deriving an analytical expression for the time derivatives of Eq. (8) to obtain an expression that relates the slack variable to the system input.
@@
$\hskip 28.45274pt+\frac{1}{2}\sum_{j=1}^{i}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t)\text{ for }i\in\{1,\dots,\rho_{k}-1\}$
$\frac{d^{\rho_{k}}}{dt^{\rho_{k}}}\left(\phi_{k}(t,x(t))+\frac{z^{2}_{k}(t)}{2}\right)=L_{g}L_{f}^{\rho_{k}-1}\phi_{k}(t,x(t))u(t)$ (10)
$\text{ }+L_{f}^{\rho_{k}}\phi_{k}(t,x(t))+z_{k}(t)z_{k}^{(\rho_{k})}(t)+\frac{1}{2}\sum_{j=1}^{\rho_{k}}{\rho_{k}\choose j}z_{k}^{(\rho_{k}-j)}(t)z_{k}^{(j)}(t).$
@@
By General Leibniz rule [37, p. 318], $\frac{d^{n}u(x)v(x)}{dx^{n}}=\sum_{k=0}^{n}\binom{n}{k}\frac{d^{k}u(x)}{dx^{k}}\frac{d^{n-k}v(x)}{dx^{n-k}}$, it follows that the $i\geq 0$ derivatives of $\frac{1}{2}z_{k}(t)^{2}$ satisfy
$\frac{d^{i}}{dt^{i}}(\frac{1}{2}z_{k}(t)^{2})$ $=\frac{1}{2}\sum_{j=0}^{i}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t)$
$=z_{k}(t)z_{k}^{(i)}(t)+\frac{1}{2}\sum_{j=1}^{i-1}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t).$
@@
is inevitable for all $(t,x)\in[t_{0},\infty)\times\mathbb{R}^{n}$. Note, $r=m$ (see Eq. (7)) implying $\Omega_{g}$ is a square matrix.
@@
where $\tilde{z}:=[z_{1},...,z_{1}^{(\rho_{1}-1)},...,z_{r},...,z_{r}^{(\rho_{r}-1)}]^{\top}\in\mathbb{R}^{\sum_{k=1}^{r}\rho_{k}}$ and $n_{A}=n+\sum_{k=1}^{r}\rho_{k}$.
@@
$f_{A}(t,x_{A})$ $:=\begin{bmatrix}f(t,x)-g(t,x)\Omega_{g}^{-1}(t,x)\Omega_{f}(t,x)\\ [z_{1}^{(1)},\dots,z_{1}^{(\rho_{1}-1)}]^{\top}\\ 0\\ \vdots\\ [z_{r}^{(1)},\dots,z_{r}^{(\rho_{r}-1)}]^{\top}\\ 0\end{bmatrix}\in\mathbb{R}^{n_{A}},$ (15)
$g_{A}(t,x_{A})$ $:=\begin{bmatrix}-g(t,x)\Omega_{g}^{-1}(t,x)D(z)\\ 0_{(\rho_{1}-1)\times r}\\ \mathcal{I}_{r,1}\\ \vdots\\ 0_{(\rho_{r}-1)\times r}\\ \mathcal{I}_{r,r}\end{bmatrix}\in\mathbb{R}^{n_{A}\times r}.$
@@
By Assumption 1 it follows $\phi$ has relative degree $\rho$ (Definition 1). Given an integrable input, $w\in L^{1}([t_{0},T],\mathbb{R}^{r})$, let $z_{i}^{(\rho_{i})}(t):=w_{i}(t)$. Since the virtual input $w$ is integrable, it follows by the fundamental theorem of calculus that each of the anti-derivatives of $w$ is absolutely continuous (see Corollary 3.33 in [38, p. 105]). Hence, for each $1\leq i\leq\rho_{j}$ there exists an absolutely continuous function $z_{i}^{(j-1)}(t):=\int_{t_{0}}^{t}z_{i}^{(j)}(s)ds+z_{i}^{(j-1)}(t_{0})$ almost everywhere. Define $u$ according to Eq. (13). By rearranging Eq. (13), we get Eq. (11). From Lemma 1 it then follows

### v2

@@ Abstract
Feedback Linearisation (FBL) is a widely used technique that applies feedback laws to transform input-affine nonlinear control systems into linear control systems, allowing for the use of linear controller design methods such as pole placement. However, for problems with state constraints, controlling the linear system induced by FBL can be more challenging than controlling the original system. This is because simple state constraints in the original nonlinear system become complex nonlinear constraints in the FBL induced linearised system, thereby diminishing the advantages of linearisation. To avoid increasing the complexity of state constraints under FBL, this paper introduces a method to first augment system dynamics to capture state constraints before applying FBL. We show that our proposed augmentation method leads to ill-defined relative degrees at state constraint boundaries. However, we show that ill-defined relative degrees can be overcome by using a switching FBL controller. Numerical experiments illustrate the capabilities of this method for handling state constraints within the FBL framework.
@@ 1 Introduction
Nonlinear systems pervade every facet of our world, from turbulent flows [1] to the sophisticated operations of human-made cybernetic infrastructure [2]. Unlike linear systems, nonlinear systems do not obey the solution superposition property leading to phenomena such as non-global stability, limit cycles and sensitivity to small initial changes (chaos). These properties make the control of nonlinear systems significantly more challenging than that of linear systems. Nevertheless, there exist several techniques that approximate or transform challenging nonlinear problems into simpler linear problems, such as Jacobian/Taylor linearisation, Koopman operators [3], Carleman linearisation [4] and Feedback Linearisation. FBL is a fairly mature topic based on the early works [5, 6] from the seventies and has been extensively treated in textbooks [7].
@@
The primary contribution of this paper is to propose a general FBL method capable of handling state constraints without the need for any online optimisation. The following three aspects summarise the specific contributions of the paper:
  • Scalable Constraint Handling: We extend the work of [29] to cases where the constraint dimension is greater than the input dimension by introducing the novel method of sequentially capturing the constraints into system dynamics, and then an unconstrained augmented system is derived. For the first time, an analytic expression for the transformation in [29] is derived and shown in Lemma 1.
  • Fundamental Controller Design Principles: Our main result, Theorem 1, establishes the fundamental principles for designing implementable controllers that obey state constraints. This theorem reveals that the properties of integrability and invertibility of virtual controllers are key when certifying constraint satisfaction under our state augmentation approach.
    Firstly, we develop an integral controller framework in Section 3.1 to ensure a bounded controller. Next, we resolve an inherent limitation of the transformation technique in [29], more specifically, in Proposition 1, we theoretically prove that the transformation may lead to ill-defined relative degrees at the boundary of the constraints. Therefore, in Sections 5.1 and 5.2, a controller switching strategy [30] [31] is introduced to overcome the invertibility challenge caused by ill-defined relative degrees.
  • Establishing a Practical Implementation Workflow: Moving beyond existing results in the literature [30] [31], we provide a definitive controller implementation method in Section 5.3 that bridges the gap between theory and practice, in which back-substitution is employed to extinguish the dependency on augmented states, thereby enabling a switchable FBL controller to be implemented on the original system.
@@
Furthermore, our method is summarised in the flowchart given in Fig. 1, follows three steps: 1) Augment the dynamics to capture state constraints and modify with integral control structure, 2) Apply FBL to the integral augmented system to derive an FBL controller, 3) Implement the FBL controller where augmented state dependencies are removed via algebraic substitution.
@@
Extending the state-space by augmenting the system dynamics has previously been used to solve discrete-time problems with non-additively separable cost functions [32, 33]. Our work focuses on tracking problems in continuous time rather than minimising general cost functions, using state augmentation to encapsulate state constraints. We follow and extend the approach of [29] by introducing a slack variable to represent the square root of the proximity to the state constraint boundary. Analogous to the first step of FBL, we differentiate this equation to establish a relationship between the slack variable and the system input. Solving for the input, we obtain a new higher-dimensional system where the virtual input corresponds to the highest derivative of the slack variable. We show that the augmented system will not violate state constraints for any integrable input. We overcome the limitation of [29], that the number of input dimensions must match the number of constraints required, by proposing a sequential approach where we iteratively augment the state space to capture constraints. We also differ from [29] by imposing an integral controller structure to ensure any input used is integrable and hence cannot violate the state constraints. Moreover, unlike in [29], which uses HJB equations to solve for the controller, we synthesise a closed-loop controller using FBL and analyse the associated challenges that arise from this approach.
@@
Unfortunately, it is non-trivial to apply FBL out of the box to the resulting augmented system, since we show that the augmented system has an ill-defined relative degree, changing its value along the state constraint boundary. Therefore, in order to apply FBL, we use the switching controller from [30] [31] that was designed to handle cases of ill-defined relative degree. Finally, we synthesise a state constraint feasible closed-loop controller for the original system coupled with integral states by removing the augmented states such that the controller only depends on the states of the original system and the induced integral states. In Fig. 2, we implement the proposed method on a Single-Input Single-Output (SISO) system studied in [29], and stabilise the state trajectory to the origin while guaranteeing that the state constraint is satisfied, in contrast to the classical FBL method, where the constraints are violated. We show more implementation details in Example 1 and manually derive the controller following our proposed method.
Figure 2: Figure associated with Example 1 showing the trajectory deduced by our proposed method against that of the classical FBL method.
@@
While our underlying methodology is applicable to MIMO systems, the implementation involving a switching controller is restricted to SISO systems. Accordingly, Sections 2 to 4, which extend the state augmentation approach of [29] and analyse the complications that arise when combining this method with FBL, are presented in the general MIMO setting. From Section 5 onwards, however, we focus exclusively on SISO systems, where the practical implementation of the FBL controller, particularly the switching logic, is more tractable.
@@
We denote the set of $k$-differentiable functions, $f:X\to Y$ by $C^{k}(X,Y)$, with $C^{\infty}(X,Y)$ denoting the set of infinitely differentiable functions. Along the same lines, we denote $L^{1}(X,Y)$ as the set of Lebesgue integrable functions. Given $F\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{r})$ and $x\in C^{1}([0,\infty),\mathbb{R}^{n})$ we define $\frac{dF(t,x(t))}{dt}=\frac{\partial F(t,x(t))}{\partial x}\frac{\partial x(t)}{\partial t}+\frac{\partial F(t,x(t))}{\partial t}$, where $\frac{\partial F(t,x)}{\partial x}=[\frac{\partial F(t,x)}{\partial x_{1}},\dots,\frac{\partial F(t,x)}{\partial x_{n}}]\in\mathbb{R}^{1\times n}$ is a row vector. We define the 0-th derivative as $\frac{d^{0}}{dt^{0}}F(t,x)=F(t,x)$ and $\frac{\partial^{0}}{\partial t^{0}}F(t,x)=F(t,x)$. Consider $\psi\in C^{i}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{m})$, $\psi^{(i)}(t,x(t))$ denotes the i-th order time derivative of $\psi(t,x(t))$. Given $f\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{n})$, $g\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{n\times m})$ and $h\in C^{1}([0,\infty)\times\mathbb{R}^{n},\mathbb{R}^{r})$, for $i\in\{1,\dots,r\}$, we denote the Lie derivatives of the i-th component in $h$ by $L_{f}^{k}h_{i}\in\mathbb{R}$ and $L_{g_{j}}L_{f}^{k}h_{i}\in\mathbb{R}$ that are given in Eqs. (1),
$L_{f}^{0}h_{i}(t,x)=h_{i}(t,x),$
$L_{f}^{k}h_{i}(t,x)=\frac{\partial L_{f}^{k-1}h_{i}(t,x)}{\partial x}f(t,x)+\frac{\partial L_{f}^{k-1}h_{i}(t,x)}{\partial t}\text{ for }k\in\mathbb{N},$
$L_{g_{j}}L_{f}^{j}h_{i}(t,x)=\frac{\partial L_{f}^{k}h_{i}(t,x)}{\partial x}g_{j}(t,x)\text{ for }k\in\mathbb{N}\cup\{0\},$ (2)
where $g_{j}$ denotes the j-th column in $g$ for $j\in\{1,\dots,m\}$. Then, we define a collection vector $L_{g}L_{f}^{k}h_{i}(t,x):=[L_{g_{1}}L_{f}^{k}h_{i}(t,x),\dots,L_{g_{m}}L_{f}^{k}h_{i}(t,x)]\in\mathbb{R}^{1\times m}$. This definition of Lie derivative is useful when dealing with time-varying systems, see [34] for more details.
@@ 2 Problem Formulation:
ODE (4) is commonly expanded as $\dot{x}(t)=f(t,x(t))+\sum_{i=1}^{m}g_{i}(t,x(t))u_{i}(t)$ [35] to demonstrate the correspondence between the input matrix and control vector fields, where $g_{i}\in\mathbb{R}^{n}$ denotes the i-th column in $g$.
@@ 3 State Augmentation: From
The slack variable, $Z$, depends explicitly on the state of the system through Eq. (8) and thus implicitly depends on the input of the system. By computing the explicit expression that relates the system input to the slack variable, we may substitute the input variable for the slack variable, treating the derivative of the slack variable as a pseudo-input variable of the transformed system. Then, the transformed system will remain state-feasible for any value taken by the pseudo-input, since by the quadratic nature of the slack variable it follows from Eq. (8) that $\phi(t,x(t))\geq 0$ for any $t\in[t_{0},\infty)$. The slack variables $z_{i}$ for all $i\in\{1,\dots,r\}$ are augmented to form new state-space coordinates for the transformation system. In the work of [29], the time derivatives of Eq. (8) were not analytically derived. Next, we extend this work by deriving an analytical expression for the time derivatives of Eq. (8) to obtain an expression that relates the slack variable to the system input.
@@
$\qquad\qquad\qquad\qquad\qquad\quad+\frac{1}{2}\sum_{j=1}^{i}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t)\text{ for }i\in\{1,\dots,\rho_{k}-1\}$
$\frac{d^{\rho_{k}}}{dt^{\rho_{k}}}\left(\phi_{k}(t,x(t))+\frac{z^{2}_{k}(t)}{2}\right)=L_{g}L_{f}^{\rho_{k}-1}\phi_{k}(t,x(t))u(t)+L_{f}^{\rho_{k}}\phi_{k}(t,x(t))$ (10)
 $\qquad\qquad\qquad\qquad\qquad\quad\text{ }+z_{k}(t)z_{k}^{(\rho_{k})}(t)+\frac{1}{2}\sum_{j=1}^{\rho_{k}}{\rho_{k}\choose j}z_{k}^{(\rho_{k}-j)}(t)z_{k}^{(j)}(t).$
@@
By General Leibniz rule [38, p. 318], $\frac{d^{n}u(x)v(x)}{dx^{n}}=\sum_{k=0}^{n}\binom{n}{k}\frac{d^{k}u(x)}{dx^{k}}\frac{d^{n-k}v(x)}{dx^{n-k}}$, it follows that the $i\geq 0$ derivatives of $\frac{1}{2}z_{k}(t)^{2}$ satisfy
$\frac{d^{i}}{dt^{i}}(\frac{1}{2}z_{k}(t)^{2})=$ $\frac{1}{2}\sum_{j=0}^{i}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t)=z_{k}(t)z_{k}^{(i)}(t)$
$+\frac{1}{2}\sum_{j=1}^{i-1}{i\choose j}z_{k}^{(i-j)}(t)z_{k}^{(j)}(t).$
@@
is invertible for all $(t,x)\in[t_{0},\infty)\times\mathbb{R}^{n}$. Note, $r=m$ (see Eq. (7)) implying $\Omega_{g}$ is a square matrix.
@@
where $\tilde{z}:=[z_{1},\dots,z_{1}^{(\rho_{1}-1)},\dots,z_{r},\dots,z_{r}^{(\rho_{r}-1)}]^{\top}\in\mathbb{R}^{\sum_{k=1}^{r}\rho_{k}}$ and $n_{A}=n+\sum_{k=1}^{r}\rho_{k}$.
@@
$f_{A}(t,x_{A}):=\begin{bmatrix}f(t,x)-g(t,x)\Omega_{g}^{-1}(t,x)\Omega_{f}(t,x)\\ [z_{1}^{(1)},\dots,z_{1}^{(\rho_{1}-1)}]^{\top}\\ 0\\ \vdots\\ [z_{r}^{(1)},\dots,z_{r}^{(\rho_{r}-1)}]^{\top}\\ 0\end{bmatrix}\in\mathbb{R}^{n_{A}},$
$g_{A}(t,x_{A}):=\begin{bmatrix}-g(t,x)\Omega_{g}^{-1}(t,x)D(z)\\ 0_{(\rho_{1}-1)\times r}\\ \mathcal{I}_{r,1}\\ \vdots\\ 0_{(\rho_{r}-1)\times r}\\ \mathcal{I}_{r,r}\end{bmatrix}\in\mathbb{R}^{n_{A}\times r}.$ (15)
@@
By Assumption 1 it follows $\phi$ has relative degree $\rho$ (Definition 1). Given an integrable input, $w\in L^{1}([t_{0},T],\mathbb{R}^{r})$, let $z_{i}^{(\rho_{i})}(t):=w_{i}(t)$. Since the virtual input $w$ is integrable, it follows by the fundamental theorem of calculus that each of the anti-derivatives of $w$ is absolutely continuous (see Corollary 3.33 in [39, p. 105]). Hence, for each $1\leq i\leq\rho_{j}$ there exists an absolutely continuous function $z_{i}^{(j-1)}(t):=\int_{t_{0}}^{t}z_{i}^{(j)}(s)ds+z_{i}^{(j-1)}(t_{0})$ almost everywhere. Define $u$ according to Eq. (13). By rearranging Eq. (13), we get Eq. (11). From Lemma 1 it then follows
@@
Note, Theorem 1 applies to the original system (4) after the augmented states are eliminated via back-substitution using Eq. (16).